OpenStudy (anonymous):

how can i factor: a^(4n)-b^(4n) Seems to me its already in its simplest form

OpenStudy (anonymous):

Factor and simplify are different things.

OpenStudy (epikrebel):

(a-b)^4n

OpenStudy (anonymous):

\[a^{4n}-b^{4n}\]\[=(a^{2n})^{2}-(b^{2n})^{2}=...\]

OpenStudy (epikrebel):

If they share something, it can be factored.

OpenStudy (anonymous):

I'm sure if you expand (a-b)^4n, you won't get a^4n - b^4n

OpenStudy (anonymous):

okay yeah that was a stupid question :P I understand it now! thanks anyways guys!

OpenStudy (anonymous):

\[\large{\color{blue}{a^{4n}-b^{4n} = (a^{2n})^2 - (b^{2n})^2 = (a^{2n} + b^{2n} )(a^{2n}-b^{2n})}}\]

OpenStudy (anonymous):

a^2n - b^2n can be further factorised!

OpenStudy (anonymous):

The identity used here is \(\large{\color{red}{a^2-b^2=(a+b)(a-b)}}\)

OpenStudy (anonymous):

\[\large{(a^{2n}+b^{2n})(a^{2n}-b^{2n}) = (a^{2n}+b^{2n})[(a^n)^2 - (b^n)^2] \\= (a^{2n}+b^{2n}) (a^n + b^n)(a^n-b^n)}\]

OpenStudy (anonymous):

but a^(2n)+b^(2n) cant be furthur factorized?

OpenStudy (anonymous):

Yes @tok1997 a^(2n)+b^(2n) can not be further factorised but a^(2n)-b^(2n) can be.

OpenStudy (anonymous):

Yup! unless... using complex number...

OpenStudy (anonymous):

yeah ... @RolyPoly

OpenStudy (anonymous):

thanks guys :)

OpenStudy (anonymous):

You're welcome @tok1997